OpenStudy (giz05):

Robert weighs 144 pounds one month before he plans to run a marathon. If he loses 3 percent of his weight while training, what is his weight on the day of the marathon?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Recall: $$\Huge x\text{%}=\dfrac{a}{b}$$

OpenStudy (ribhu):

144- 4.32 139.68 pounds

OpenStudy (skullpatrol):

If he loses 3% of his weight, then he must weigh 100% — 3% = 97% of 144 pounds, right?

OpenStudy (giz05):

3% of Robert's weight = (0.03)(144 lbs) = 4.3 lbs If Robert loses 3% of his weight while training, on the day of the marathon he will weigh 144 - 4.3 = 139.7 lbs which is approximately 140 lbs this is the best i got . I hope I'm right ...

OpenStudy (skullpatrol):

97% of 144 $$\dfrac{97}{100}•144$$

OpenStudy (skullpatrol):

Yes, you are right :-) Check it with 0.97•144